Use De Moivre’s theorem to simplify the following
(a) (cos π/5+i sin π/5)10, (b) (cos π/9+i sin π/9)−3, (c) {cos(−π/6) + i sin(−π/6)}−4
1
Expert's answer
2020-05-25T20:28:50-0400

Use De Moivre’s theorem: "(\cos x+i \sin x)^n=\cos(nx )+i \sin(nx)"


(a) "(\cos \pi/5+i \sin \pi/5)^{10}=\cos(10\times \pi/5 )+i \sin(10\times \pi/5)=\cos (2\pi)+i \sin(2\pi)=1"


(b) "(\cos \pi/9+i \sin \pi/9)^{-3}=\cos(-3\times \pi/9 )+i \sin(-3\times \pi/9)=\cos (-\pi/3)+i \sin(-\pi/3)=\frac{1}{2}-i \frac{\sqrt 3}{2}"


(c) "(\cos (-\pi/6)+i \sin (-\pi/6))^{-4}=\cos(-4\times (-\pi/6) )+i \sin(-4\times (-\pi/6))=\cos (2\pi/3)+i \sin(2\pi/3)=-\frac{1}{2}+i \frac{\sqrt 3}{2}"